hey, i have two questions i really need some help on. i'm beyond lost and really need an explanation on how to get the answers. thanks ahead for your time.(:

here are the problems and directions;;

1.] write an equation for the line that is parallel to the give line and that passes through the given point.

y = -7/2x + 6 ; (-4,-6)

2.] find the slope of a line perpendicular to the graph of the equation.

y = -x/5 - 7

Answer by oscargut(2103)

write an equation for the line that is parallel to the give line and that passes through the given point.
y = -7/2x + 6 ; (-4,-6)
the new line has the same slope then y=-7/2x+b
and passes through (-4,-6) then -6=-(7/2)(-4)+b
then -6=14+b then b=-20
answer: y=-(7/2)x-20
find the slope of a line perpendicular to the graph of the equation.
y = -x/5 - 7
if m and n are slopes of perpendicular lines then m*n=-1
here m =-1/5 then n =5
answer: slope is 5